Reports 1-1 of 1 Clear search Modify search
Online Data Analysis (Omega)
verkindt - 11:10 Wednesday 13 July 2011 (29894) Print this report
The 3 kHz glitches have disappeared
Good news: since the relock of July 8th 16:00 UTC, the glitches around 3 kHz have disappeared from the Omega plots (see plots 1 and 2).
Not so good news: I have not found any logbook entry that could be associated to this change. Only TCS tuning (entry 29860) or Bas entries (29861 and 29862) are around this time.
Images attached to this report
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×